Suppose that market demand for a good is given by Q = 9 - 0.3 P while market supply is given by Q = 4 + 0.3 P where Q is the quantity of the good in units, and P is the price of the good in $ per unit. Calculate the consumer surplus in this market if the market is perfectly competitive (Don't use the $ sign in your answer and round it to two decimal points)
Answer #1

Answer

The market is in equilibrium at Qd=Qs

9-0.3P=4+0.3P

0.6P=5

P=8.33

Q=9-0.3*8.33

Q=6.5

-----------

inverting demand curve to inverse form to find Y-intercept of the demand curve

Q=9-0.3P

0.3P=9-Q

P=9/.3 -(1/0.3)*Q

P=30-(1/0.3)*Q

the Y-axis intercept is 30

----------

CS=0.5*(Y-intercept of the demand curve -P)*Q

=0.5*(30-8.33333)*6.5

=$70.4166775

the consumer surplus is 70.42